OpenStudy (anonymous):

Need help. I just need it to be checked. I'm stuck between two answers and i just need to know which one right. it will be below

OpenStudy (anonymous):

\[\frac{ 6a^2-13a-5 }{ 4a-1 }\times \frac{ 8a^2-18a-5 }{ 3a^2+a }\]

OpenStudy (anonymous):

thats the question^^^

OpenStudy (anonymous):

and these are the two answers I'm stuck between. A: \[\frac{ (4a+1)(2a-5)^2 }{ a(4a-1) }\] B: \[\frac{ (4a+1)(2a-5) }{a(4a-1) }\]

OpenStudy (dumbcow):

well after factoring there should be 4 factors on top, 3 on bottom if the "3a+1" cancels, that leaves 3 factors on top, 2 on bottom ...

OpenStudy (anonymous):

But i dont have a choice that has 3 factors on top and two on the bottom.

OpenStudy (dumbcow):

hmm where do you think the (2a-5)^2 came from? --> (2a-5)(2a-5)

OpenStudy (anonymous):

ohh, hahaha. I wasnt really thinking. so it would be A right? @dumbcow

OpenStudy (dumbcow):

yep :)

OpenStudy (anonymous):

okay thank you! :)